# Provably Overwhelming Transformer Models with Designed Inputs

### v1

@@ 1 Introduction
Empirical Results: In section 6, we use our algorithm in practice on a trained, single-layer model. The algorithm checks for “overwhelming” when restricting the free string to an element of one of two permutation classes. We use fixed strings of varying sizes (up to 500 tokens): our fixed strings come from either: (1) the test set of our training data, (2) randomly sampled tokens, (3) a repeated string. We also test whether “overwhelming” persists across multiple steps of token generation.
@@ 2 Model
In this paper, we focus on a standard single-layer transformer model which uses RoPE positional encoding, parallel feed-forward residuals, and layer normalization. Our model closely follows the GPT-NeoX architecture (Black et al., 2022) except that: (1) we do not use an attention mask, and (2) we use a single layer-normalization for the MLP and the attention mechanism instead of two  [^2] Ref. (Black et al., 2022) intended to initially use a single layer-normalization for the MLP and attention mechanism, but, due to a bug, this was changed in the final implementation. We thus use the intended architecture in this paper..
@@
$\mathcal{M}=\mathrm{Unembed}\circ(\texttt{Attn}+\text{MLP}+\mathbb{I})\circ\texttt{LN}\circ\text{E}$
@@ 3 Worst-Case Deviation and a Custom Norm
It will be convenient to “lift” the input space into two parts and then find bounds on the worst-case deviation by optimizing over the two parts separately.
Definition 3.3 (Lifted Worst-Case Deviation).
For a function $f:\mathcal{X}\to\mathbb{R}^{n}$ where $\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}$ and $f$ is well defined on $\mathcal{Y}$ and $\mathcal{Z}$, we define the lifted worst-case deviation of $f$ as
$\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})_{p}:=\sup_{Y_{1},Y_{2}\in\mathcal{Y}}\sup_{Z_{1},Z_{2}\in\mathcal{Z}}\norm{f(Y_{1},Z_{1})-f(Y_{2},Z_{2})}_{p}$
In appendix C, we state and prove some useful properties of the worst-case deviation. Mainly, we note that $\mathcal{W}$ behaves mostly like a norm but has the nice property that “lifting” the domain is monotonic. I.e. for lifted domain $\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}$, $\mathcal{W}(f;\mathcal{X})\leq\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})$.
@@ 4.1 Algorithm for deciding Overwhelming
  • bound the worst-case deviation of each component as a function of the blowup and shift from layer-normalization.
@@
  Calculate upper-bound $W^{f^{enc}}$ as in lemma 4.15.
  Calculate
$W=\mathrm{Lip}(\mathrm{Unembed})$
$\cdot\mathopen{}\mathclose{{}\left(W^{\texttt{Attn}}+\mathrm{Lip}({\text{MLP}})_{\infty}\cdot W^{f^{enc}}+W^{f^{enc}}}\right)$
as per lemma 4.14 and
@@
Discussing normalization as a function of expectation and variance is a bit cumbersome. So, we will first define an analagous “blowup” and “shift” for normalization.
Definition 4.6 (Blowup and Shift Sets).
For some $X\in\mathbb{O}_{des}$, Let $B_{j}(X)$ (for blowup) denote
$\frac{1}{\sqrt{\mathrm{Var}[X\text{E}\cdot e_{j}^{T}]+\epsilon}}\cdot\gamma$
and let $S_{j}(X)$ (for shift) denote
$\frac{-\mathbb{E}[X\text{E}\cdot e_{j}^{T}]}{\sqrt{\mathrm{Var}[X\text{E}\cdot e_{j}^{T}]+\epsilon}}\cdot\gamma+\beta.$
We define $\mathcal{B}_{j}=\{B_{j}(X)|X\in\mathbb{O}_{des}\}$, and $\mathcal{S}_{j}=\{S_{j}(X)|X\in\mathbb{O}_{des}\}$. Moreover, for ease of notation, we will define
=$\mathcal{B}=\{(B_{1}(X),\dots,B_{d_{emb}}(X)\mid X\in\mathbb{O}_{des}\}$
$\mathcal{S}=\{(S_{1}(X),\dots,S_{d_{emb}}(X))\mid X\in\mathbb{O}_{des}\}$
and the set $\mathcal{B}\mathcal{S}$ to equal
$\mathopen{}\mathclose{{}\left\{\bigg{(}(B_{1}(X),\dots,B_{d_{emb}}(X)),(S_{1}(X),\dots,S_{d_{emb}}(X))\bigg{)}}\right\}$
where $X\in\mathbb{O}_{des}$. Finally, let $B^{\min}_{j}=\min(\mathcal{B}_{j})$, $B^{\max}_{j}=\max(\mathcal{B}_{j})$, $S^{\min}_{j}=\min(\mathcal{S}_{j})$ and $S^{\max}_{j}=\max(\mathcal{S}_{j})$.
We will now break down the model in a way such that it is easier to analyze. For component ${\texttt{comp}}\in\{\texttt{Attn},\text{MLP},\mathbb{I}\}$, we will use $f^{{\texttt{comp}}}:\mathbb{O}^{n_{ctx}}\times\mathcal{B}\mathcal{S}\to\mathbb{R}^{{n_{ctx}}\times{d_{vocab}}}$ to denote the following function:
$f^{{\texttt{comp}}}(X,(B,S))={\texttt{comp}}\circ f^{enc}(X,(B,S))$
where
$f^{enc}(X,(B,S))=(X\cdot\text{E}\cdot\mathrm{diag}(B)+S).$
Note that $f^{enc}(X,(B(X),S(X)))$ is exactly equal to $\texttt{LN}(X\cdot\text{E})$ for layernorm function LN.
@@
$\mathcal{M}=\mathrm{Unembed}\circ(f^{\texttt{Attn}}+f^{\text{MLP}}+f^{\mathbb{I}})$
and so, by the monotonicity of lifting for worst-case deviation,
@@
$\leq$ $\sum_{{\texttt{comp}}}\mathcal{W}(\mathrm{Unembed}\cdot\vec{e}_{n_{ctx}}\cdot f^{{\texttt{comp}}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})$
@@
4.2.1 Bounding Blowup and Shift
We can get rather straightforward bounds on the blowup and shift of the model.
Lemma 4.7 (Blowup and Shift Bounds).
We bound blowup and shift: for every $B_{j}\in\mathcal{B}_{j}$ and $S_{j}\in\mathcal{S}_{j}$
$\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\max}+\epsilon}}\leq B_{j}\leq\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\min}+\epsilon}}$
and
$\min\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)\leq S_{j}$
$\leq\max\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)$
where $\mu^{\min}_{j},\mu^{\max}_{j}$ are the lower and upper bounds on the expectation of the $j$-th column of $X\cdot\text{E}$. $\mathrm{Var}_{j}^{\min},\mathrm{Var}_{j}^{\max}$ are similarly defined for the variance. Both are bounded in lemma D.1 within appendix D.
Moreover, we let $B^{\min}$ be a vector of the minimum values of the blowup and $B^{\max}$ be a vector of the maximum values of the blowup. Define $S^{\min}$ and $S^{\max}$ similarly for the shift.
@@ 4.2.2 Bounding Attention
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\texttt{Attn}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq\max_{h}\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}_{h}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
@@
Algorithm 4 computes the minimum and maximum values of the logits $\ell$ given point-wise bounds $B^{\min}$, $B^{\max}$, $S^{\min}$, and $S^{\max}$ as specified in lemma 4.7.
@@
In the following, $E[[n_{fix}]\cup\{{n_{ctx}}\},:]\in\mathbb{R}^{n_{fix}\times{d_{emb}}}$ denotes the matrix with the rows in $[n_{fix}]\cup\{{n_{ctx}}\}$ selected.
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot$
$\max_{B,S}\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot\mathrm{diag}(B)+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(E\cdot\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}$
$+\norm{V}_{\infty}\cdot{\beta}_{free}^{\max}\cdot\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}\}))_{\infty}$
@@
Building on the above, we obtain a worst-case deviation bound using point-wise upper and lower bounds on $B\in\mathcal{B}$ and $S\in\mathcal{S}$.
Corollary 4.13.
Let $B^{\min}$, $B^{\max}$, $S^{\min}$, and $S^{\max}$ be as defined at the end of Definition 4.6. Then, we have
$\mathcal{W}$ $(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq$
$2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\bigg{(}\max_{B\in[B^{\min},B^{\max}]}\|E[[n_{fix}]\cup\{{n_{ctx}}\},:]$
$\cdot\mathrm{diag}(B)\|_{{F_{\infty}}}+\max\big{(}\|{S^{\min}}\|_{\infty},\|{S^{\max}}\|_{\infty}\big{)}\bigg{)}$
$+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(EB+S)\cdot V}_{{F_{\infty}}}$
$+\norm{V}_{\infty}\cdot{\beta}_{free}^{\max}\cdot\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}))_{\infty}$
where the maximization can be calculated with a simple linear program 333 We use the notation $B\in[B^{\min},B^{\max}]$ to denote the set of vectors, $B$, where $B_{j}^{\min}\leq B_{j}\leq B_{j}^{\max}$. .
@@ 4.2.3 Bounding MLP and Identity
Because the MLP and identity components are simple feed-forward networks without any “cross-talk” between tokens, we can use simple Lipschitz bounds.
@@
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\text{MLP}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq\mathrm{Lip}({\text{MLP}})_{\infty}\cdot\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
@@
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\mathbb{I}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$=\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}.$
4.2.4 Bounding the Encoding Function
We upper bound the worst-case deviation for $f^{enc}$ using bounds on the “blowup and shift” in lemma 4.7:
Lemma 4.15 (Worst-case deviation of $f^{enc}$).
$\mathcal{W}$ $(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq$
$\max_{t\in[{d_{vocab}}]}$ $\max_{j\in[{d_{emb}}]}\max_{B\in[\vec{0},B^{\max}-B^{\min}]}$
$\mathopen{}\mathclose{{}\left|X[t]\cdot\text{E}\cdot\mathrm{diag}(B)\cdot\vec{e}_{j}}\right|+S_{j}^{\max}-S_{j}^{\min}$
where the inner maximum term can be computed via a simple linear program.
@@ 5 Overwhelming Under Permutation Invariance
In section 4, we provided a concrete algorithm that decides overwhelming for a fixed context size ${n_{ctx}}$. In this section, we provide a more refined algorithm which can provide a tighter bound on the worst-case deviation in the restricted setting where the free tokens are allowed to be any permutation of a fixed string. Up until this point, we have been “lifting” the domain of $\mathbb{O}_{des}$ into $\mathbb{O}_{des}\times\mathcal{B}\mathcal{S}$ to prove worst-case deviation bounds: i.e. we have been separating out the effect of layer normalization when proving a fixed string to be overwhelming.
In this section, we nullify the effect of layer normalization by defining an equivalence relation on $\mathbb{O}_{des}$ where the blowup and shift sets are the same for all elements in the equivalence class. Then, we can derive a simpler upper-bound on the worst-case deviation of the model over the equivalence class as we can treat the layer normalization as a constant factor!
@@
To bound the worst-case deviation over an equivalence class we use a similar algorithm (algorithm 2) to algorithm 1. We still bound the worst-case deviation of attention, but now the blowup and shift sets are singletons, leading to much tighter bounds. Additionally, since blowup and shift are constant, the worst-case deviations of $f^{\text{MLP}},f^{\mathbb{I}}$ are 00, and, we can use a linear program to find the extremal softmax contributions for the free tokens.
@@
  Set $B=\frac{\gamma}{\sqrt{\mathrm{Var}[X\cdot E]+\epsilon}}$
  Set $S=-\mathbb{E}[X\cdot E]B+\beta$
  Set $f^{enc}(\vec{e})=\vec{e}\cdot E\cdot\mathrm{diag}(B)+S$
@@
  Calculate $W=\mathrm{Lip}(\mathrm{Unembed})\cdot W^{\texttt{Attn}}$
@@ 5.1 Bounding Attention
Worst-case deviation of attention is bounded as follows for fixed blowup $B$ and shift $S$ when considering the permutation class $[X]$ with free tokens FreeToks:
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$
$\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot B+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\norm{(E[(n_{fix},{n_{ctx}}),:]B+S)\cdot V}_{{F_{\infty}}}.$
The proof follows analogously to that of lemma 4.12 except for two main differences: (1) we have a further restriction on the free tokens and thus only need to consider $E[(n_{fix},{n_{ctx}}),:]$ instead of $E$ and (2) becuase the blowup and shift are fixed, $\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S})_{\infty}=0$ and so wecan drop the last term in the original bound.

### v2

@@ 1 Introduction
Empirical Results: In section 6, we use our algorithm in practice on a trained, single-layer model. The algorithm checks for “overwhelming” when restricting the free string to an element of one of two permutation classes. We use fixed strings of varying sizes (up to 1,500 tokens): our fixed strings come from either: (1) the test set of our training data, (2) randomly sampled tokens, (3) a repeated string. We also test whether “overwhelming” persists across multiple steps of token generation.
@@ 2 Model
In this paper, we focus on a standard decoder-only transformer model which uses RoPE positional encoding, parallel feed-forward residuals, and layer normalization. Our model closely follows the GPT-NeoX architecture (Black et al., 2022) except that we use a single layer-normalization for the MLP and the attention mechanism instead of two  [^2] Ref. (Black et al., 2022) intended to initially use a single layer-normalization for the MLP and attention mechanism, but, due to a bug, this was changed in the final implementation. We thus use the intended architecture in this paper..
@@
$\mathcal{M}$ $=\mathrm{Unembed}\circ\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\texttt{Attn}^{(i)}+\text{MLP}^{(i)}}\right)\circ\texttt{LN}+\mathbb{I}}\right)\circ\text{E}$
@@ 3 Worst-Case Deviation and a Custom Norm
Then, we can define the worst-case deviation as the follows.
@@
In appendix C, we state and prove some useful properties of the worst-case deviation.
@@ 4.1 Algorithm for deciding Overwhelming
  Set $W=W^{\texttt{Attn}}$ as per lemma 4.12 and
@@
In a single-layer model, layer-norm turns out to be quite simple and almost trivial. Simply put, we can take the layer-norm function and re-write it as a change of basis from the basis specified by the embedding matrix, E, to a basis specified by a “normalized” embedding matrix.
Definition 4.6 (Layernorm Matrix, $\text{E}_{\texttt{LN}}$).
For row $i\in[{d_{vocab}}]$ of the embedding matrix, $E[i]\in\mathbb{R}^{d_{emb}}$, define the normalized vector $\vec{n}_{i}$ as
$\vec{n}_{i}=\texttt{LN}(E[i])=\frac{E[i]-\mathbb{E}[E[i]]\cdot\vec{1}}{\sqrt{\mathrm{Var}[E[i]]+\epsilon}}\cdot\gamma+\beta\cdot\vec{1}.$
Then, define the matrix $\text{E}_{\texttt{LN}}\in\mathbb{R}^{{d_{vocab}}\times{d_{emb}}}$ as $\text{E}_{\texttt{LN}}[i]=\vec{n}_{i}$.
We will now break down the model in a way such that it is easier to analyze. For component ${\texttt{comp}}\in\{\texttt{Attn},\text{MLP},\mathbb{I}\}$, we will use $f^{{\texttt{comp}}}:\mathbb{O}^{n_{ctx}}\to\mathbb{R}^{{n_{ctx}}\times{d_{vocab}}}$ to denote the following functions:
$f^{\mathbb{I}}(X)$ $=\mathrm{Unembed}\circ\mathbb{I}\circ E\cdot X,$
$f^{\text{MLP}}(X)$ $=\mathrm{Unembed}\circ\text{MLP}\circ\text{E}_{\texttt{LN}}\cdot X,$
$f^{\texttt{Attn}}(X)$ $=\mathrm{Unembed}\circ\texttt{Attn}\circ\text{E}_{\texttt{LN}}\cdot X.$
@@
$\mathcal{M}$ $=f^{\texttt{Attn}}+f^{\text{MLP}}+f^{\mathbb{I}}$
and so, by the triangle inequality for worst-case deviation,
@@
$\leq$ $\sum_{{\texttt{comp}}}\mathcal{W}(\cdot\vec{e}_{n_{ctx}}\cdot f^{{\texttt{comp}}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})$
@@ 4.2.1 Bounding Attention
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\texttt{Attn}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}$
$\leq\max_{h}\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}_{h}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}.$
@@
Algorithm 4 computes the minimum and maximum values of the logits $\ell$.
@@
In the following, let $\text{E}_{\texttt{LN}}[[n_{fix}]\cup\{{n_{ctx}}\},:]\in\mathbb{R}^{n_{fix}\times{d_{emb}}}$ denotes the matrix with the rows in $[n_{fix}]\cup\{{n_{ctx}}\}$ selected.
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des})_{\infty}\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})$
$\cdot\norm{\text{E}_{\texttt{LN}}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot V\cdot\mathrm{Unembed}}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\norm{\text{E}_{\texttt{LN}}\cdot V\cdot\mathrm{Unembed}}_{{F_{\infty}}}.$
@@ 4.2.2 Bounding MLP and Identity
Because the MLP and identity components are simple feed-forward networks without any “cross-talk” between tokens and the last token of the model is fixed within our scheme, we can trivially see that the worst-case deviation of these components are 00.
@@
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\text{MLP}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}=0$
@@
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{\mathbb{I}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}=0.$
4.2.3 Causal Masking in Attention
We note that a causal mask will not change the attention scores for the last token in a sequence (i.e. the query token). And so, the attention scores for the query token will be the same regardless of whether we use a causal mask or not. Our outlined algorithm (algorithm 1) thus still provides a valid upper bound on the worst-case deviation for models with and without causal masking.
@@ 5 Overwhelming Under Permutation Invariance
In section 4, we provided a concrete algorithm that decides overwhelming for a fixed context size ${n_{ctx}}$. In this section, we provide a more refined algorithm which can provide a tighter bound on the worst-case deviation in the restricted setting where the free tokens are allowed to be any permutation of a fixed string. We introduce a new technique to achieve tighter bounds: specifically, we use an integer-linear program relaxation to model the “global” constraint on the free tokens.
@@
To bound the worst-case deviation over an equivalence class we use a similar algorithm (algorithm 2) to algorithm 1. We still bound the worst-case deviation of attention, but now the set of free tokens are restricted, leading to a tighter bound.
@@
  Set $f^{enc}(\vec{e})=\vec{e}\cdot\text{E}_{\texttt{LN}}$
@@
  Set $W=W^{\texttt{Attn}}$
@@ 5.1 Bounding Attention
Worst-case deviation of attention is bounded as follows when considering the permutation class $[X]$ with free tokens freeToks:
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des})_{\infty}$
$\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\|\text{E}_{\texttt{LN}}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot V$
$\cdot\mathrm{Unembed}\|_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\norm{\text{E}_{\texttt{LN}}[(n_{fix},{n_{ctx}}),:]\cdot V\cdot\mathrm{Unembed}}_{{F_{\infty}}}.$
The proof follows analogously to that of lemma 4.11 except for two main differences: (1) we do not necessarily use all of the free tokens in the dictionary and can thus restrict out search, and (2) we have some global structure on the free tokens and thus can use a linear program to find the extremal softmax contributions.